OpenStudy (anonymous):

an automobile traveling 95 km/h overtakes a 1.00km long train travelling in the same direction on a track parallel to the road. if the train's speed is 75km/h, how long does it take the car to pass it and how far will the car have traveled in this time? what are the results if the car and train are traveling in opposite directions?

OpenStudy (irishboy123):

relative velocity = 95 - 75 = 20km/h relative distance = 1km time = 1/20 h = 3mins = 180s in other direction rel vel = 95 + 75 = 170km/h time = 1/170 h = 21.2s